### v1

@@ Abstract
Current methods to extract relational triples directly make a prediction based on a possible entity pair in a raw sentence without depending on entity recognition. The task suffers from a serious semantic overlapping problem, in which several relation triples may share one or two entities in a sentence. It is weak to learn discriminative semantic features relevant to a relation triple. In this paper, based on a two-dimensional sentence representation, a bi-consolidating model is proposed to address this problem by simultaneously reinforcing the local and global semantic features relevant to a relation triple. This model consists of a local consolidation component and a global consolidation component. The first component uses a pixel difference convolution to enhance semantic information of a possible triple representation from adjacent regions and mitigate noise in neighbouring neighbours. The second component strengthens the triple representation based a channel attention and a spatial attention, which has the advantage to learn remote semantic dependencies in a sentence. They are helpful to improve the performance of both entity identification and relation type classification in relation triple extraction. After evaluated on several publish datasets, it achieves competitive performance. Analytical experiments demonstrate the effectiveness of our model for relational triple extraction and give motivation for other natural language processing tasks.
@@ 1 Introduction
Relational triple extraction is defined as a task to identify pairs of entities and extract their relations in the form of (subject, relation, object) or (s, r, o) from unstructured text. Instead of identifying relations between predefined entity types, the relational triple extraction task tries to find semantic relationships between any entity pairs in a sentence. It has the advantage to find all possible entity relations in a sentence and avoid the cascading failure caused by predicting entity types. In recent years, the task has achieved great attention in natural language processing and is widely adopted to support other downstream tasks such as knowledge graph construction (Dong et al., 2014), information retrieval (Sinha et al., 2015), etc.
Techniques to support relational triple extraction can be roughly divided into two frameworks: pipeline framework and joint framework. In pipeline framework (Chan and Roth, 2011; Zhong and Chen, 2021), the task is implemented as two independent subtasks: entity recognition and relation extraction. The main problem is that they suffer from error propagation and ignore the interaction between two subtasks. Recent researches focus on joint methods, which extract the triple (s, r, o) in an end-to-end multi-objective framework. In addition, to avoid the error propagation, joint models also have the ability to share model parameters between different tasks. They showed great potential to support relational triple extraction (Zheng et al., 2017; Wang et al., 2020; Ren et al., 2021; Yu et al., 2020).
Even great successes have been achieved in this field, relational triple extraction is still a challenging task, because a sentence usually contains several overlapped entity pairs. The overlapping phenomenon can be divided into four categories: normal, entity pair overlapping (EPO), single entity overlapping (SEO) and subject object overlapping (SOO) (Zheng et al., 2021). The normal refers to relational triples without overlapping. EPO denotes to two triples containing the same two entities but annotated with different relation types. In SEO, two triples share a same subject or object entity. In SSO, both subject and object are simultaneously shared by two relational triples. An example of the overlapping phenomenon is shown in Figure 1.
@@
The above sentence contains three relational triples. The “Live in” relation and “Locate in” relation share the same entity “Palo Alto”, which leads to the SEP overlap. On the other hand, the sentence also contains a SOO overlap between the “Locate in” relation and “Work for” relation, where the entity “Tesla” acts as the subject and object of the two relational triples, respectively. Because overlapped relational triples share the same contextual features, it is difficult to distinguish them, especially in a sequential sentence representation, where a relational triple is usually represented by fusing two entity representations.
Recent works have a trend to transform a sentence into a two-dimensional (2D) sentence representation, e.g., the planarized sentence representation (Geng et al., 2023) or the table filling (Wang et al., 2020; Ren et al., 2021; Zhang et al., 2017). In the 2D representation, each element denotes an entity pair representation in a sentence. The 2D sentence representation unfolds a semantic plane. It is effective to represent all entity pairs in a sentence. Under this representation, the relational triple extraction is conducted as element classification. The main problem for relational triple extraction in the 2D representation is that entity semantics are spread out in the 2D representation, which is influential on the extraction performance. Before to show the motivation of our model, we first give a 2D sentence representation in Figure 1.
@@
Refer to caption        $(a)$ Theory Representation
Refer to caption        $(b)$ Real Representation
Figure 2 contains two 2D representations about the sentence in Figure 1, where deep colors represent the strength of entity information in an element (also known as “span”). Sub-figure (a) is a theory sentence representation. Sub-figure (b) is a real sentence representation (b), which was generated by feeding the sentence into a well trained model.
In Figure 2, the theory representation and real representation shows the same semantic patterns, which indicate four important phenomena in relational triple extraction. First, the semantic plane is usually constructed by concatenating two word representations. Adjacent elements denote to overlapped phrases which share the same contextual features. Second, two entities of a relational triple can be separated by great distances, e.g., the triple ( “Franz”, “Live in”, “Palo Alto”). Third, in the 2D representation, the semantic information of a real entity is shared by all elements in the same column or the same row, because these elements are generated from the entity representation. Fourth, a sentence often contains several relation triples which may share one or two entities. The task suffers from a serious semantic overlapping problem. Under this condition, to advance the discriminability of of deep networks, it is important to make full use of local features relevant to a named entity and to encode long-distance semantic dependences (or global features) relevant to a named entity pair.
@@
The rest of this paper is organized as follows. Section 2 provides an introduction to related works on joint entity and relation extraction. In Section 3, we present our proposed model in detail. The effectiveness of our model for joint entity and relation extraction is demonstrated through experiments conducted in Section 4. Furthermore, in Section 5, we delve into the role and performance of the various modules within the model. Finally, Section 6 offers a conclusion and outlines future works pertaining to the proposed model.
@@ 2.1 Tagging based Methods
Tagging based methods convert the process of triple extraction into a series of interdependent sequence labeling tasks. For example, Wei et al. (2020) have proposed a CasRel model, which detects all potential head entities and subsequently applies relation-specific sequence taggers to identify their corresponding tail entities. Zhan et al. (2022) presented a comprehensive approach, known as DropRel, to tag the triple boundaries and relation types via three distinct dependence relations. Furthermore, to ensure the preservation of semantic information while generating distinct vectors for a given token pair, they proposed a dropout-normalization layer. Recently, Wang et al. (2023) proposed a relation-specific sequence tagging decoder that employs a filter attention mechanism to highlight more informative features, alongside an auxiliary matrix that amalgamates information related to entity pairs. However, it is worth noting that methods based on sequence tagging often suffer from issues related to semantic loss and error propagation.
@@ 2.2 Seq2Seq based Methods
Seq2Seq based methods treat a triple as a sequence of tokens and utilize an encoder-decoder framework to produce relational triples. This process is similar to machine translation. For example, CopyRE (Zeng et al., 2018) utilized as a copy mechanism to generate the relationship between two corresponding entities. The problem is that it fails to extract the order of relational facts within a sentence. In order to address this limitation, Zeng et al. (2019) introduced a reinforcement learning to a sequence-to-sequence model. Moreover, the CopyRE model is only capable of predicting the last word of an entity. In order to overcome this constraint, CopyMTL (Zeng et al., 2020) employed as a multi-task learning framework to address the multi-token entity problem. TDEER (Li et al., 2021) proposed a novel translating decoding schema that represents the relation as a translation operation from subjects to objects. However, the Seq2Seq-based approach poses challenges in constructing a vector with rich semantic features, particularly for long sentences.
@@ 2.3 Table Filling based Methods
Table filling based methods formulate the task of triple extraction as a table filling process, in which a table is comprised of the Cartesian product of an input sentence. For example, SPTree (Miwa and Bansal, 2016) effectively captured both word sequence and dependency tree substructure information. This is accomplished through the use of stacked bidirectional tree-structured LSTM-RNNs on top of bidirectional sequential LSTM-RNNs. TF-MTRNN (Gupta et al., 2016) proposed a novel context-aware joint entity and word-level relation extraction approach through semantic composition of words. This model is capable of modeling multiple relation instances without knowing the corresponding relation arguments in a sentence. RS-TTS (Zhang et al., 2023a) predicted potential relations to avoid computational redundancy and utilizes an efficient tagging and scoring strategy for entity decoding.
In addition to the above mentioned methods, researchers have investigated other approaches. For example, Bekoulis et al. (2018) presented a Multi-head model, which formulates the relational triple extraction task as a multi-head selection problem. Li et al. (2019) implemented the task as a multi-turn question answering task, where a Multi-turn QA model was proposed. Fu et al. (2019) employed a graph convolutional network-based method known as Graphrel. Eberts and Ulges (2020) utilized a span extraction approach. Sun et al. (2021) proposed a multitask learning-based RTE model. In the field of attention, Xu et al. (2015) proposed a visual attention method to model the importance of features in the image caption task. A residual attention network (Wang et al., 2017) was proposed with a spatial attention mechanism using downsampling and upsampling. Besides, a channel attention mechanism was proposed by SENet (Hu et al., 2018).
@@ 3 Methodology
The architecture of our model is demonstrated in Figure 3. It comprises three modules: Sentence Encoding Module, Bi-consolidating Module and Triple Generating Module.
@@ 3.1 Sentence Encoding Module
Given a sentence consisting of $N$ words, it can be denoted as $\mathbf{S}=[x_{1},x_{2},\ldots,x_{N}]$ or $\mathbf{S}=x_{[1:N]}$. To capture the semantic information of each word, a commonly adopted approach is to utilize a pretrained language model that maps each word into an abstract representation. In our experiments, each sentence is fed into the BERT (Devlin et al., 2019) for learning the high-order abstract semantic features of the words:
@@
To learn structural features of a sentence, a position embedding and an attention embedding are introduced to generate a position matrix($\mathbf{M}^{P}$) and an attention matrix ($\mathbf{M}^{A}$). The purpose of position embedding is to capture the positional information or order of tokens within a sentence. On the other hand, attention embedding is derived from the raw input using an attention layer. This embedding is learned by assigning weights to input features based on their relevance or importance to the output. The two embedding are represented as follows:
@@
where $P_{ij}=N-i$ if $j\succ i$, else $P_{ij}=j-N$. Therefore, $\mathbf{M}^{P}$ and $\mathbf{M}^{A}$ are a $\mathbb{R}^{N\times N\times D_{p}}$ matrix and $\mathbb{R}^{N\times N\times D_{a}}$ matrix, respectively.
@@ 3.2.1 Local Consolidation
The local consolidation module consists of a stack of multiple convolution blocks. A convolution block comprises two components: pixel difference convolution and layer normalization (Ba et al., 2016). The convolution module is constructed by stacking ${L}$ identical convolution blocks. The default value of ${L}$ is four in our experiments. The pixel difference convolution and layer normalization layer are represented as follows:
Pixel Difference Convolution: the process of pixel difference convolution (PDC) is represented as Figure 5, where the convolution kernels operate on the local feature map patch and replace the original pixels with pixel differences.
@@
Various strategies can be employed to select pixel pairs from the representation $\mathbf{M}^{so}$. In our study, we adopt three kinds of pixel difference convolution in Su et al. (2021), named as: central PDC (CPDC), angular PDC (APDC) and radial PDC (RPDC). We also propose five new strategies to encode pixel relations from different directions (XY-axis, diagonal, omnidirectional and bidirectional) to improve the ability for capturing the local dependencies between neighbouring elements. The strategies to select pixel pairs are as illustrated in Figure 4, in which each patch denotes to a convolution kernel. An arrow between two elements represent a pixel pair from $(z_{i},z_{i}^{\prime})$.
Figure 4: From expanded LBP descriptors, eight examples of pixel difference convolution were created.
@@
Layer Normalization: when the number of convolution blocks increases during the optimization process, the gradient descent method will change the distribution of input features. To ensure the stability of these feature distributions, a layer normalization layer is employed to normalize the output of the pixel difference convolution. This normalization process enables a larger learning rate and accelerates the convergence of the model. Moreover, it is helpful to mitigate the overfitting problem, making the training process smoother.
@@ 3.2.2 Global Consolidation
The global consolidation component is introduced in this section, aiming to facilitate the model’s capture of global information from a 2D sentence representation.
As shown in Figure 3, the global consolidation component comprises a channel attention module and a spatial attention module. Given an input sentence, the channel and spatial attentions carry out complementary attention, highlighting “what” and “where”, respectively. The channel attention is applied globally, whereas the spatial attention operates on a local scale.
@@
It can be concluded from the above equation that the previous application of channel attention solely focused on the lowest frequency component of the first term GAP (Global Average Pooling). It ignores the subsequent components expressed in Equation 6, resulting in disregarded information. However, the excessive utilization of frequency components leads to the generation of superfluous information. It is imperative to take into account the optimal selection of frequency components. Therefore, in our work, we adopt an average and a max pooling operation to improve the performance.
The averaging pooling serves to eliminate redundant information while concurrently retaining the most significant information within a sentence to the greatest degree possible. In comparison, max pooling enables the identification and focuses on the most salient features presented within a given sentence. Spatial information from feature maps is aggregated through the average and max pooling operations. It generates two distinct spatial context descriptors: $\mathbf{U}^{\mathbf{c}}_{\mathbf{avg}}$, representing the average-pooled features, and $\mathbf{U}^{\mathbf{c}}_{\mathbf{max}}$, representing the max-pooled features.
@@
The representation ($\mathbf{M}^{so}_{i,j}$) is fed into the local consolidation component and the global consolidation component to learn local and global semantic features. The process is represented as:

### v2

@@ Abstract
Current methods to extract relational triples directly make a prediction based on a possible entity pair in a raw sentence without depending on entity recognition. The task suffers from a serious semantic overlapping problem, in which several relation triples may share one or two entities in a sentence. In this paper, based on a two-dimensional sentence representation, a bi-consolidating model is proposed to address this problem by simultaneously reinforcing the local and global semantic features relevant to a relation triple. This model consists of a local consolidation component and a global consolidation component. The first component uses a pixel difference convolution to enhance semantic information of a possible triple representation from adjacent regions and mitigate noise in neighbouring neighbours. The second component strengthens the triple representation based a channel attention and a spatial attention, which has the advantage to learn remote semantic dependencies in a sentence. They are helpful to improve the performance of both entity identification and relation type classification in relation triple extraction. After evaluated on several publish datasets, the bi-consolidating model achieves competitive performance. Analytical experiments demonstrate the effectiveness of our model for relational triple extraction and give motivation for other natural language processing tasks.
@@ 1 Introduction
Relational triple extraction is defined as a task to identify pairs of entities and extract their relations from unstructured text. The output can be represented in the form of triples: (subject, relation, object) or (s, r, o). Instead of identifying relations from predefined entity types, the relational triple extraction task can find relationships between any entity pairs in a sentence. It has the advantage to find all possible entity relations in a sentence and avoid the cascading failure caused by predicting entity types. In recent years, the task has achieved great attention in natural language processing and is widely adopted to support other downstream tasks such as knowledge graph construction (Dong et al., 2014), information retrieval (Sinha et al., 2015), etc.
Techniques to support relational triple extraction can be roughly divided into two frameworks: pipeline framework and joint framework. In pipeline framework (Chan and Roth, 2011; Zhong and Chen, 2021), the task is divided into two independent subtasks: entity recognition and relation extraction. The main shortcoming is that they suffer from the error propagation problem and ignore the interaction between two subtasks. Recent researches focus on joint framework, which extract the triple (s, r, o) in end-to-end multi-objective models. In addition, to avoid the error propagation, joint models also have the ability to share model parameters between different tasks. They have shown a great potential to support the relational triple extraction task (Zheng et al., 2017; Wang et al., 2020; Ren et al., 2021; Yu et al., 2020).
Even great successes have been achieved in this field, relational triple extraction is still a challenging task, because a sentence usually contains several overlapped entity pairs. The overlapping phenomenon can be divided into four categories: Normal, entity pair overlapping (EPO), single entity overlapping (SEO) and subject object overlapping (SOO) (Zheng et al., 2021). The Normal refers to relational triples without overlapping. EPO denotes to two triples containing the same two entities but annotated with different relation types. In SEO, two triples share the same subject or object entity. In SOO, both subject and object are simultaneously shared by two relational triples. An example of the overlapping phenomenon is shown in Figure 1.
@@
The above sentence contains three relational triples. The “Live in” relation and “Locate in” relation share the same entity “Palo Alto”, which leads to the SEO overlap. On the other hand, the sentence also contains a SOO overlap between the “Locate in” relation and “Work for” relation. The entity “Tesla” acts as the subject and object in the two relational triples. Because overlapped relational triples share the same contextual features, it is difficult to distinguish them, especially in a sequential sentence representation, where a relational triple is usually represented by fusing two entity representations.
Recent works have a trend to transform a sentence into a two-dimensional (2D) sentence representation, e.g., the planarized sentence representation (Geng et al., 2023) or the table filling (Li et al., 2022). In the 2D representation, each element denotes an entity pair representation in a sentence. Because the 2D sentence representation unfolds a semantic plane, it is effective to represent all entity pairs in a sentence. Based on the 2D representation, the relational triple extraction is conducted as element classification. The main problem is that entity semantics are spread out in the 2D representation, which is influential on the relational triple extraction. Before to show the motivation of our model, we first give a 2D sentence representation in Figure 1.
@@
       $(a)$ Theory Representation
       $(b)$ Real Representation
Figure 2 contains two 2D representations about the same sentence in Figure 1, where deep colors represent the strength of the entity semantic in an element (also known as “span”). Sub-figure (a) is a theory sentence representation. It indicates that, in the 2D sentence representation, the semantics of entities (e.g., “Franz”, “Tesla” and “Alto”) are shared by all elements in the same columns and rows. Sub-figure (b) is a real sentence representation. To generate this representation, the sentence is first fed into a well trained model, which was conducted with the settings in Section 4.1. Then, the real representation is visualized by using the 2D sentence representation in the trained model.
In Figure 2, the theory and real representation show the same semantic pattern, which indicates four important phenomena in relational triple extraction. First, the semantic plane is usually constructed by concatenating two word representations. Adjacent elements denote to overlapped phrases which share the same contextual features. Second, two entities of a relational triple can be separated by great distances, e.g., the triple ( “Franz”, “Live in”, “Palo Alto”). Third, in the 2D representation, the semantic information of a real entity is shared by all elements in the same column or the same row, because these elements are generated from the entity representation. Fourth, a sentence often contains several relation triples which may share one or two entities. Therefore, the relation triple extraction task suffers from serious semantic overlapping problems in the 2D sentence representation.
@@
The rest of this paper is organized as follows. Section 2 provides an introduction to related works on joint entity and relation extraction. In Section 3, the bi-consolidating model is presented and in details. The effectiveness of our model for joint entity and relation extraction is demonstrated through experiments conducted in Section 4. Furthermore, in Section 5, different components are evaluated to show the effectiveness of the proposed models in relation triple extraction. Finally, Section 6 offers a conclusion and outlines future works.
@@ 2.1 Tagging based Methods
Tagging based methods convert the process of triple extraction into a series of interdependent sequence labeling tasks. For example, Wei et al. (2020) have proposed a CasRel model, which detects all potential head entities and subsequently applies a relation-specific sequence tagger to identify their corresponding tail entities. Zhan et al. (2022) presented a comprehensive approach, known as DropRel, to tag the triple boundaries and relation types via three distinct dependence relations. Furthermore, to ensure the preservation of semantic information while generating distinct vectors for a given token pair, they proposed a dropout-normalization layer. Recently, Wang et al. (2023) proposed a relation-specific sequence tagging decoder that employs a filter-attention mechanism to highlight more informative features. However, it is worth noting that methods based on sequence tagging often suffer from issues related to semantic loss and error propagation.
@@ 2.2 Seq2Seq based Methods
Seq2Seq based methods treat a triple as a sequence of tokens and utilize an encoder-decoder framework to produce relational triples. For example, CopyRE (Zeng et al., 2018) utilized as a copy mechanism to generate the relationship between two corresponding entities. The main problem with this model is that it is weak to extract the order of relational facts within a sentence. In order to address this limitation, Zeng et al. (2019) introduced a reinforcement learning to improve this CopyRE model. However, it is only capable of predicting the last word of an entity. In order to overcome this constraint, CopyMTL (Zeng et al., 2020) employed as a multi-task learning framework to address the multi-token entity problem. TDEER (Li et al., 2021) proposed a novel translating decoding schema that represents the relation as a translation operation from subjects to objects. Even the Seq2Seq-based approaches have great potential to support relation extraction, it still poses challenges in constructing a vector with rich semantic features, particularly for long sentences.
@@ 2.3 Table Filling based Methods
Table filling based methods formulate the task of triple extraction as a table filling process. For example, SPTree (Miwa and Bansal, 2016) effectively captured both word sequence and dependency tree substructure information. This is accomplished through the use of stacked bidirectional tree-structured LSTM-RNNs on top of bidirectional sequential LSTM-RNNs. TF-MTRNN (Gupta et al., 2016) proposed a novel context-aware joint entity and word-level relation extraction approach through semantic composition of words. This model is capable of modeling multiple relation instances without knowing the corresponding relation arguments in a sentence. RS-TTS (Zhang et al., 2023a) predicted potential relations to avoid computational redundancy and utilizes an efficient tagging and scoring strategy for entity decoding.
In addition to the above mentioned methods, researchers have investigated other approaches. For example, Bekoulis et al. (2018) presented a multi-head model, which formulates the relational triple extraction task as a multi-head selection problem. Li et al. (2019) implemented the task as a multi-turn question answering task, where a multi-turn QA model was proposed. Fu et al. (2019) employed a graph convolutional network-based method known as Graphrel. Eberts and Ulges (2020) utilized a span extraction approach. Sun et al. (2021) proposed a multi-task learning-based RTE model. In the field of attention, Xu et al. (2015) proposed a visual attention method to model the importance of features in the image caption task. A residual attention network (Wang et al., 2017) was proposed with a spatial attention mechanism using downsampling and upsampling. Besides, a channel attention mechanism was proposed by SENet (Hu et al., 2018).
@@ 3 Methodology
The architecture of the bi-consolidating model is demonstrated in Figure 3. It comprises three modules: Sentence Encoding Module, Bi-consolidating Module and Triple Generating Module.
@@ 3.1 Sentence Encoding Module
Given a sentence consisting of $N$ words, it can be denoted as $\mathbf{S}=[x_{1},x_{2},\ldots,x_{N}]$ or $\mathbf{S}=x_{[1:N]}$. To capture the semantic information of each word, a common approach is to utilize a pretrained language model that maps each word into an abstract representation. In our experiments, each sentence is fed into the BERT (Devlin et al., 2019) for learning the high-order abstract semantic representations of words. The process can be formalized as:
@@
To learn structural features of a sentence, following Geng et al. (2023) and Li et al. (2022), the position embedding and attention embedding are applied to generate a position matrix($\mathbf{M}^{P}$) and an attention matrix ($\mathbf{M}^{A}$) encoding semantic dependencies between elements. The two embedding are represented as follows:
@@
where $P_{ij}=N-i$ if $j>i$, else $P_{ij}=j-N$. Therefore, $\mathbf{M}^{P}$ and $\mathbf{M}^{A}$ are a $\mathbb{R}^{N\times N\times D_{p}}$ matrix and $\mathbb{R}^{N\times N\times D_{a}}$ matrix, respectively.
@@ 3.2.1 Local Consolidation
The local consolidation component consists of a stack of multiple convolution blocks. A convolution block comprises two layers: a pixel difference convolution layer and a normalization layer (Ba et al., 2016). The local consolidation component is constructed by stacking ${L}$ identical convolution blocks. The default value of ${L}$ is four in our experiments. The pixel difference convolution and layer normalization layer are introduced as follows:
Pixel Difference Convolution: the process of pixel difference convolution (PDC) is represented in Equation (5), where the convolution kernels operate on a local patch and replace the original pixels with pixel differences.
@@
Various strategies can be employed to select pixel pairs from the representation $\mathbf{M}^{so}$. In our study, we adopt three kinds of pixel difference convolution presented in Su et al. (2021), named as: central PDC (CPDC), angular PDC (APDC) and radial PDC (RPDC). We also propose five new strategies to encode pixel relations from different directions (XY-axis, diagonal, omnidirectional and bidirectional) to improve the ability for capturing the local dependencies between neighbouring elements. The strategies to select pixel pairs are as illustrated in Figure 4, in which each patch denotes to a convolution kernel. The arrows between elements represent a pixel pair from $(z_{i},z_{i}^{\prime})$, which indicate the direction of semantic dependencies between them.
Figure 4: Eight strategies for selecting pixel pairs are created.
@@
Layer Normalization: when the number of convolution block is increased during the optimization process, the gradient descent method will change the distribution of input features. To ensure the stability of these feature distributions, a layer normalization layer is employed to normalize the output of the pixel difference convolution. This normalization process enables a larger learning rate and accelerates the convergence of the model. Moreover, it is helpful to mitigate the overfitting problem, making the training process smoother.
@@ 3.2.2 Global Consolidation
The global consolidation component is set to facilitate the model’s capture of global information from a 2D sentence representation. As shown in Figure 3, the global consolidation component comprises a channel attention module and a spatial attention module. The channel attention is applied globally, whereas the spatial attention operates on a local scale. Therefore, channel attention and spatial attention are complementary to each other. Spatial attention highlights keywords related to entities or relations in a given sentence, while channel attention captures global semantic features between entities and relations.
@@
It can be concluded from the above equation that the previous application of channel attention solely focused on the lowest frequency component of the first term GAP (Global Average Pooling). It ignores the subsequent components expressed in Equation (6), resulting in disregarded information. However, the excessive utilization of frequency components leads to the generation of superfluous information. It is imperative to take into account the optimal selection of frequency components. Therefore, in our work, we adopt an average and a max pooling operation to improve the performance.
The averaging pooling serves to eliminate redundant information while concurrently retaining the most significant information within a sentence. In comparison, max pooling enables the identification and focuses on the most salient features presented within a sentence. Spatial information from feature maps is aggregated through the average and max pooling operations. It generates two distinct spatial context descriptors: $\mathbf{U}^{\mathbf{c}}_{\mathbf{avg}}$, representing the average-pooled features, and $\mathbf{U}^{\mathbf{c}}_{\mathbf{max}}$, representing the max-pooled features.
@@
The representation ($\mathbf{M}^{so}_{i,j}$) is simultaneously fed into the local consolidation component and the global consolidation component to learn local and global semantic features. The process is represented as: